2. Find the area of a quadrant of a circle whose circumference is 22 cm.

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
Circumference of the circle, C = 22 cm (given) It should be noted that a quadrant of a circle is a sector which is making an angle of 90°. Let the radius of the circle = r As C = 2πr = 22, R = 22/2π cm = 7/2 cm ∴ Area of the quadrant = (θ/360°) × πr2 Here, θ = 90° So, A = (90°/360°) × π r2 cm2 = (49/16) π cm2 = 77/8 cm2 = 9.6 cm2